Sam has a certain number of rabbits and birds inside a cage. When Sam counted all the animals, there were 35. When he counted all their legs, there were 94. How many rabbits were there in Sam's cage? ( answer is 12 - show your solution )
Answer by macston(5194)

B=no of birds; R= no of rabbits
B+R=35 (No.of animals)
2B+4R=94 (no. of legs)
Solve first equation for R and substitute in second
R=35-B
2B+4(35-B)=94
2B+140-4B=94 Subtract 140 from each side
-2B+140-140=94-140
-2B=-46 divide each side by -2
B=23 ANSWER ! There are 23 birds
B+R=35
23+R=35 Subtract 23 from each side
R=12 ANSWER 2 There are 12 rabbits
CHECK:
23+12=35 total animals
2(23)+4(12)=94
46+48=94
94=94 Total legs